Hot News

 


   Return to Phoenix Home Page


This page notes recent updates, news which might not yet be incorporated into the manuals or problems encountered during observing runs which will (hopefully) be repaired in the near future. We intend to update this frequently.

13 August 2015: The final Phoenix run at the Mayall concluded on 8 June 2015, with Caty Pilachowski gathering the final photons in a successful observing run. Phoenix has been shipped downtown and is being prepared for transport to Gemini South, where it is intended to be scheduled as a visitor instrument. The interface for the Gemini telescope has been shipped from Chile to Tucson, cleaned up and repainted, and installed on Phoenix. The dark slide motor (unused at Kitt Peak but necessary at Gemini) has been installed and initial tests indicate that all of the motors and the ThAr line emission lamp operate correctly under computer control. The current vacuum port actuator, which is somewhat tricky to operate, will be replaced by a more conventional vacuum valve. Since there is no plan to use Phoenix on Kitt Peak in the future, it will be shipped to Chile with spare computers, motors, and everything required for future use. Stay tuned for progress reports on this.

16 April 2015: In response to requirements for an upcoming observing run, we have removed the J8265 filter (which was never used) in the Filter Wheel with a new J9440 filter. These changes have been documented on the Phoenix Filter page and in the latest version (2.04) of the 4-m Phoenix manual. The decommissioning of the 2.1-m telescope has made the manual for that telescope irrelevant.

16 May 2013: Both the 2.1-m and 4-m manuals have been updated to version 2.0, dated 15/16 May 2013. The startup procedures have been further simplified, some minor display nuances have been corrected, and an Engineering version of the Phoenix Daemon window now opens on the initial startup for low-level diagnostic purposes. A separate Phoenix Saver window has been created for logging of the saved images. Because of the generous display real estate now available with current monitors, the use of a single window for multiple tasks is no longer necessary.

Additional technical manuals describing the setup and cooldown of Phoenix at the 2.1-m and low-level commands and procedures have been added to the Technical Manuals page.

27 February 2013: For observer startups: The command newobserver will prompt for the OBSERVER and PROPID fields which will appear in the FITS headers. This allows one to enter new observers without having to run the dreaded obsinit routine. However, this change works only as long as the current WildFire session. If one exits WildFire, it may be necessary to either do a full WildFire startup (Start Wildfire rather than Start Phoenix), or execute newobserver after restarting WildFire.

21 February 2013: The User Manuals for the 2.1-m and 4-m have been revised again, to V1.03, dated 21 February 2013. The updated manuals include the notes from 19 February regarding autodisplay on the ds9 and the revised grating setting offset. A new script ab has also been included in the available Phoenix scripts; it is similar to abba, except that it carries out only one integration on the two slit positions.

We are also requesting that on the final night of an observing block, that the observer or the instrument change crew move the viewer wheel to the dark position with the command viewer dark. This will shield the detector from potential contamination which can occur when the instrument is warmed up or when it is cooled down for the succeeding observing session.

19 February 2013: When one initially brings up the Wildfire program, images may not automatically display on the ds9. If this is the case, change to the data directory in the IRAF window and manually display an image using "display image.fits 1". If the image displays correctly in the ds9, then succeeding Phoenix images should autodisplay in the ds9.

The offset in the grating setting has changed from 1800 to -1000. Thus, in going to a desired frequency, enter "grat {frequency in cm-1} K -1000".

Currently, the collimator focus position is 236.4.

The Phoenix ITC is now working properly.

31 January 2013: The Phoenix webpages have been updated to include (along with this page), new User manuals for the 2.1-m and 4-m telescopes, as well as the traditional pages of useful information about Phoenix.

The User manuals contain information on setting the grating for a particular wavelength/frequency, but during the latest 4-m run in December 2012, we found that the zero point offset of the grating had changed significantly (3000 units) for unknown reasons. The focus setting was also significantly different from that which had been optimized during the previous 2.1-m run. It is possible that this could have happened during the transport from the 2.1-m to the 4-m, but the cause remains unknown. We plan to recalibrate both the grating and focus settings at the beginning of the February 2013 run.

The ITC pages for Phoenix provide correct information, but the plots do not work. This is evidently a result of the upgraded IRAF server not working properly, not a problem with the ITC. This problem is being addressed.



rjoyce@noao.edu
13 August 2015